OpenStudy (babynini):

Limit of artan(e^x) as x approaches 0

OpenStudy (babynini):

@Astrophysics

OpenStudy (babynini):

is it just a rule that it equals pi/2 every time? xD

OpenStudy (priyar):

isn't it pi/4 ?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \lim_{x\to0}~\arctan(e^x) }\) Yes, but it is a direct answer too.

OpenStudy (babynini):

Nein, it's pi/2

OpenStudy (solomonzelman):

pi/4 is correct, which you will notice by graphing. Also, \(\color{#000000 }{ \displaystyle \lim_{x\to0}~\arctan(e^x)=\arctan(1)=\pi/4 }\) by a direct sub.

OpenStudy (solomonzelman):

I am saying this, not to confuse with the incorrect answer.

OpenStudy (priyar):

we can check like this if in doubt: tan (pi/4)=1

OpenStudy (babynini):

hmmm

jimthompson5910 (jim_thompson5910):

tan(pi/2) is undefined because cos(pi/2) = 0 tan(x) = sin(x)/cos(x)

OpenStudy (babynini):

In class when we went over this the prof presented this by just saying lim as x approaches 0 artan(e^x) = pi/2 as if it were a rule though. Like wolframalpha also says it is pi/4 but the man said pi/2 xD

OpenStudy (babynini):

And..the homework agrees o.o -.- o.o

OpenStudy (priyar):

hey there its x tends to infinity..

OpenStudy (solomonzelman):

More evidence: http://www.wolframalpha.com/input/?i=limit+x--%3E+0+arctan(e%5Ex)

OpenStudy (priyar):

u said x tends to zero

OpenStudy (babynini):

oh crap. Sorry!

OpenStudy (solomonzelman):

Oh, infinity, indeed would change it :)

OpenStudy (babynini):

haha I was confused too because I kept getting pi/4 as well. I am so sorry xD

OpenStudy (priyar):

its ok..

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle \lim_{x\to\infty}\arctan(e^x) =\pi/2 }\) that would be right

OpenStudy (babynini):

Ok, and how do we get that it's pi/2? hehe

OpenStudy (rational):

First, you might want to review the domain and range of arctan(x)

OpenStudy (babynini):

so it's by looking at the graph and just knowing that?

OpenStudy (zenmo):

Substitute e^x as t

OpenStudy (solomonzelman):

Yes, exactly.

OpenStudy (zenmo):